OpenStudy (anonymous):

By method od delta-epsilon, Prove that lim x->1(2x^2-3x) = 5

OpenStudy (watchmath):

Is that suppose to be 2x^2+3x? because otherwise the limit is not 5!

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

no wait one second

OpenStudy (anonymous):

yes its 2x^2+3x

OpenStudy (anonymous):

hi

OpenStudy (watchmath):

Analysis: Given \(\epsilon >0\) we would like to find \(\delta >0\) such that for all \(0<|x-1|<\delta\) we have \(|2x^2+3x-5|<\epsilon\) Notice that \(\begin{align*} |2x^2+3x-5|<\epsilon \iff |2x+5||x-1|<\epsilon \end{align*}\) In order to be able to say something about |x-1| we need to bound the |2x+5| expression. If \(|x-1|<1\) then by triangle inequality \(|x|-1\leq |x-1|<1\). Hence \(|x|<2\). It follows that \(|2x+5|\leq 2|x|+5<9\). So now in order to make \(|2x+5||x-1|<\epsilon\) it is enough by making \(9|x-1|<\epsilon\), i.e. \(|x-1|<\epsilon/9\). So which $\delta$ we need to choose to make \(|x-1|<1\) and \(|x-1|<\epsilon/9\) at the same time? Choose \(\delta:=\min \{1,\epsilon/9\}\). Here is the rigorous proof (the analysis above is to come up with the following proof) Given \(\epsilon >0\) choose \(\delta:=\{1,\epsilon/9\}\). Since $\delta <1$, if $|x-1|<\delta$ then $|x-1|<1$. It follows that $|x|-1<1$. Hence $|x|<2$. If follows that \(|2x+5|\leq 2|x|+5<2(2)+5=9\). Now \[|2x^2+3x-5|=|2x+5||x-1|< 9|x-1|<9(\epsilon/9)=\epsilon\]